OpenStudy (anonymous):

-1 /2(root2)+ root3 /2(root2)

OpenStudy (anonymous):

\[\frac{-1}{\sqrt{2}}+\frac{\sqrt{3}}{\sqrt{2}}\] \[\frac{\sqrt{3}-1}{\sqrt{2}}\] \[\frac{\sqrt{3}-1}{\sqrt{2}}\cdot\frac{\sqrt{2}}{\sqrt{2}}\] \[\frac{\sqrt{6}-\sqrt{2}}{2}\]

OpenStudy (jiteshmeghwal9):

\[\Huge{\frac{-1}{2\sqrt{2}}+\frac{\sqrt{3}}{2\sqrt{2}}}\]by taking common denominators\[\Huge{\frac{-1+\sqrt{3}}{2\sqrt{2}}}\]now rationalize\[\Huge{\frac{-1+\sqrt{3}}{2\sqrt{2}}\times \frac{2\sqrt{2}}{2\sqrt{2}}}\]

OpenStudy (jiteshmeghwal9):

u will gt\[\frac{\sqrt{6}-\sqrt{2}}{2}\]